OpenStudy (anonymous):

which of these ordered pair is a solution to the inequality y-2x<=-3? a.)(3,4) b.)(1,-1) c.)(-2,3) d.)(2,4)

OpenStudy (blurbendy):

Just put x and y where they belong and see if it results in a true statement. for example the first one would be: 4 - 2(3) <= -3 4 - 6 <= -3 -2 <= -3 this is not true, however, because -2 is greater than -3

OpenStudy (anonymous):

ok thank you

OpenStudy (blurbendy):

which one is correct?

OpenStudy (anonymous):

did you figure it out yet ?